# Modeling Future Conversation Turns to Teach LLMs to Ask Clarifying Questions

### v1

@@ Abstract
Large language models (LLMs) must often respond to highly ambiguous user requests. In such cases, the LLM’s best response may be to ask a clarifying question to elicit more information. We observe existing LLMs often respond by presupposing a single interpretation of such ambiguous requests, frustrating users who intended a different interpretation. We speculate this is caused by current preference data labeling practice, where LLM responses are evaluated only on their prior contexts. To address this, we propose to assign preference labels by simulating their expected outcomes in the future turns. This allows LLMs to learn to ask clarifying questions when it can generate responses that are tailored to each user interpretation in future turns. In experiments on open-domain QA, we compare systems that trained using our proposed preference labeling methods against standard methods, which assign preferences based on only prior context. We evaluate systems based on their ability to ask clarifying questions that can recover each user’s interpretation and expected answer, and find that our training with our proposed method trains LLMs to ask clarifying questions with a 5% improvement in F1 measured against the answer set from different interpretations of each query.
@@ 1 Introduction
We propose an alternative method for annotating double-turn preferences over clarifying questions where annotators interact with the LLM by providing their responses to clarifying questions and observe the LLM’s subsequent responses. Each annotator then assigns preferences based on whether the completed interaction successfully fulfilled the request. When comparing multiple clarifying questions and direct answer responses, each annotator’s preferences are aggregated to identify the response that maximizes preference across all annotators. We depict our proposed double-turn preference annotation method and compare against standard methods in Figure 1.
To demonstrate the benefits of training LLMs to ask clarifying questions with double-turn preferences, we experiment on open-domain QA (Chen et al., 2017). We develop an automatic evaluation framework for evaluating clarifying questions that use simulated user interactions (Section 2). We evaluate systems on two axis: efficiency measured by the number of model turns and effectiveness measured by F1 score between the predicted answer set for each user and their expected gold answer.
Figure 1: Depiction of our interaction scenario and preference labeling schemes. We aim to build an LLM that can interact with user to generate the final answer set $R$, containing an answer for each user, for the input query $x$. In this example, we include two responses from state-of-the-art LLMs ([A] from GPT-4 and [B] from Gemini, full responses in Appendix A), which both presuppose an interpretation of the word football. We also include two clarifying responses ([C] and [D]) where [C] correctly disambiguates the two intended interpretations across all users. We depict two different ways to assign preferences on LLM’s initial output (Single-Turn/Double-Turn).
In our experiments, we demonstrate that training systems with double-turn preferences outperforms training with standard preferences annotation methods (Zhu et al., 2023), resulting in consistent 4-5% improvement in F1 score over three different base LLMs (Dubey et al., 2024; Mesnard et al., 2024). We also demonstrate that double-turn preferences can be used to train systems that judiciously determine if clarification is needed, or if the user’s question can be directly answered without additional interaction.
@@ 2.1 Interaction Scenario
Setting / Goal
We assume an input query $x$, which can be potentially ambiguous, and a set of $k$ users $u_{1},...u_{k}$ who posed the input query $x$. Each $i$-th user $u_{i}$ has an disambiguated intent for query ($x_{i}$) and an expected gold answer ($y_{i}$). Our goal is to build a system $M$ that returns an answer set $R$ that can satisfy as many users as possible (by providing the expected answer $y_{i}$ for each user $u_{i}$) with the minimum number of interactions (measured by the number of model turns).
We will have a simple function $\phi(M(x))$ that returns True if $M(x)$ is a clarifying question, and False otherwise. We further assume an oracle user simulation model $\psi(x,M(x),y_{i})=a_{i}$, that takes first turn input $x$, model’s initial output $M(x)$, and $i$-th user’s targeted answer $y_{i}$, that generates a response to the model’s initial response $M(x)$.
Dialogue Turns
In the initial turn, user poses the input query $x$. In the second turn, given the input query $x$, the system $M$ is to generate its initial response ($r_{\text{init}}$). Depending on whether $r_{\text{init}}$ is a question, it will trigger different dialogue trajectories.
@@
Existing open-domain QA models mostly generates an single answer to the initial input query $x$, even when the query is ambiguous. Such systems that predict an answer without first confirming the user’s intent may mislead users with incorrect answers and fail to serve the entirety of their user population. In this work, we explore methods for QA-with-Clarification, where models can ask the user a clarifying question $q$ and observe the user’s response $a$ before predicting an answer. We will also build a system that sometimes poses a clarifying question and sometimes directly answers, judiciously determine when clarifying questions is useful (Zhang & Choi, 2023a).
Evaluations
We evaluate the system in two axis: efficiency and effectiveness. To measure efficiency, we simply measure the average number of total turns from the system. For unambiguous inputs, asking clarifying question is unnecessary. If the system can satisfy users without incurring additional interaction, such system should be preferred.
To measure effectiveness, we use F1 between the model generated answer set ($R$) and the gold reference answer set ($\{y_{1},..y_{k}\})$, which provides us a holistic view of the precision and recall of our systems. When evaluating QA-with-Clarification systems, we additionally enforce that answers must be correct for their corresponding clarifying answer. We hold QA-with-Clarifications to this higher standard, as it aligns with our goal of to developing systems that can identify different user interpretations of a query and accurately predict corresponding answers.
@@ 2.2 Evaluation Framework Implementation
Data
We will leverage existing open-domain QA datasets (Kwiatkowski et al., 2019; Min et al., 2020) where each query is paired with annotated answers from multiple annotators.
Identifying clarifying question $\phi(\cdot)$
We will use a very simple method, which takes the input string and check whether it starts with “Clarifying Question:”.
User Simulation Model $\psi$ for Turn 3
We use GPT-4 as our oracle user simulation model ($\psi$). We prompt GPT-4 with the input query $x$, clarifying question $q$, and target answer $y_{i}$ and task it with generating a corresponding clarifying answer $((x,q,y_{i})\rightarrow a_{i})$. We prompt GPT-4 to abstain from providing a clarifying answer $a_{i}$ if it judges that none exists, in which case we count the resulting target answer prediction $r^{i}_{\text{next}}$ as incorrect. We further ensure that the generated clarifying answers do not leak the target answer by removing instances where the target answer $y_{i}$ appears in the predicted clarifying answer $a_{i}$, treating such examples the same as abstains from GPT-4. We provide the prompt used in Appendix B.
@@ 3 Fine-tuning LLMs to ask Clarifying Questions
We present our approach to build a model that can engage with users for a multi-turn interaction. Specifically, we focus on building a LLM $M$ which can ask clarifying question when it can satisfy more users in the future turns.
Following the standard RLHF training pipeline (Ouyang et al., 2022), we first construct an instruction tuning dataset focusing on clarifying responses as desired outputs. We then use this constructed dataset for supervised finetuning (SFT), before performing preference learning between two possible responses. We first describe our method for constructing this instruction tuning dataset before introducing our annotation scheme for assigning preferences over clarifying questions.
@@ 3.1 Preference Data Generation from User-LLM Interactions
To train LLMs to generate useful clarifying questions and identify when to ask them, we propose a novel method for annotating preferences based on full user-LLM interactions. Our method for labeling preferences deviates from standard single-turn annotation methods in two ways. First, we simulate additional interaction with the user and derive reward from the LLM’s final response after the additional turn. Second, we rely on multiple annotators for each example. While some prior work has explored using multiple annotator judgments in single-turn preference datasets (Wang et al., 2024; Kopf et al., 2023), these works aggregate annotator judgments via majority choice to generate a single preference label per example. This practice, however, can lead to adverse outcomes where systems whose outputs are catered a toward a single, majority accepted response (Fleisig et al., 2023; Santy et al., 2023). In contrast, we use preferences across multiple annotators without removing individual annotator judgments, helping us identify whether a clarifying question or direct-answer response successfully accommodates the range of different annotator’s interpretations.
@@
In our double-turn annotation scheme (depicted in Figure 2), annotators are provided an input query $x$ with several candidate clarifying question $q$ and direct-answer responses $\widehat{y}$. Annotators are tasked with responding to each clarifying question by providing the clarifying answer that best matches their own interpretation. We then show annotators the model’s final output prediction, given each clarification. Annotators then assign their preferences over clarifying questions and direct-answer responses based on whether the final output matched their expected interpretation. We then aggregate these preferences to identify which response was preferred by the most annotators. To prevent systems from asking unnecessary clarifying questions, ties between direct-answer and clarifying question responses are broken by favoring the direct-answer response. The final aggregated preferences can then be used in conjunction with standard RL methods from human preferences for training. While this annotation scheme can be applied to human annotators, in this work we use simulate user interactions for annotation.
@@
To simulate different user interpretations of the input query, we utilize datasets consisting of an input query $x$ and a set of correct answer outputs $\{y_{1},...,y_{k}\}$. For each query, we use the set of gold outputs to simulate $k$ different users, one for each answer in $\{a_{1},...,a_{k}\}$. We then simulate the behaviors of each of these users and their responses to a proposed clarifying question $q$ using a we use a trained user-simulator model, which takes as input query $x$, clarifying question $q$, and the user’s annotated answer $y_{i}$ and is trained to predict the user’s clarifying answer $(x,q,y_{i})\rightarrow a_{i}$. While we use prompted GPT-4 to simulate clarifying answers from users for evaluation, during training we use this trained user-simulator model to reduce costs and to ensure that our systems do not overfit to a particular user-simulator method.
@@ 4 Experiments
We use Llama2-7b (Touvron et al., 2023), Gemma-7b (Mesnard et al., 2024), and Llama3-8b (Dubey et al., 2024) as our base LLMs. We forego using the instruction-tuned models due to data leakage concerns, as such systems are finetuned on the entirety of NQ-Open. For training and inference, we use 8-bit quantization (Dettmers et al., 2022) with LoRA (Hu et al., 2022; Dettmers et al., 2024) (training details in Appendix B).
@@
  • AmbigQA consists of a subset of queries from NQ-Open which have been additionally annotated with whether or not the input is ambiguous. If the input is ambiguous, annotators then provide the additional answers to all possible interpretations of the query. This process identifies about 56% of all queries in NQ-Open as ambiguous, and recovers 2.6x more answers to ambiguous questions than the base NQ-Open annotations alone. We provide additional comparisons in the answer sets identified by each dataset in Appendix A.
@@
  • Answers from Model Predictions: We build feasible answer sets $Y_{\text{model}}$ to for queries in the NQ-Open training set using answer candidates proposed by a base LLM. Prior work has demonstrated that base LLMs are well calibrated in open-domain QA tasks with few-shot prompting (Kadavath et al., 2022; Cole et al., 2023). Drawing on these observations, we construct a feasible answer set $Y_{\text{model}}$ by sampling 5-shot predictions from Llama-2-7b (Touvron et al., 2023). For each query, we a sample 5-shot prompt and generate its greedy answer with temperature $T=0.0$ and sample an answer with $T=1.0$. We repeat the process 10 times per query and remove repeated answers to generate a set of model-identified feasible answers. We further filter examples where none of the feasible answers match any annotated answers from NQ-Open.
@@ 4.1 Compared Systems
Figure 3 summarizes all models used in our experiments. We have two types of models, model that will be used at the second turn (which we refer as Turn 2 model) and model that will be used at fourth turn (which we refer as Turn 4 model). The model that will be used at the second turn can ask clarification question or directly answer the question. All models that are capable of asking clarification questions are paired with Turn 4 model. We use a separate model for performing this latter step to ensure that differences in performance only reflect the value of the clarifying question, and explore training joint models for both steps in Section 5. We describe each of these below.
Direct Ans SFT
@@
This model is trained to predict the answer given the input query $(x)\rightarrow y$. We fine-tune base LLM with next-token prediction loss on the full set of question answer pairs (i.e., each $(x,y_{i})$) in our generated SFT training sets. We use this model under two inference settings: (1) Greedy where we take the single greedy answer prediction and (2) Sampled where we sample a set predicted answers from the LLM. We sample 20 answers from the model and select the top $k$ most frequent answers, where $k$ matches the gold number of annotated answers. In rare cases ($<2\%$), there are less than $k$ unique answers within the 20 samples.
Clarify SFT
@@
This model is trained to ask clarifying question to input query. We finetune base LLMs on our SFT datasets constructed with both human annotations and model predicted answers. In Section 5, we provide ablation experiments training over each dataset to compare their utility.
Clarify DPO (Starling) / Clarify DPO (Ours)
These two models are each fine-tuned from Clarify SFT model with using DPO (Rafailov et al., 2023), a method for optimizing LLMs on preference data that eliminates the need for training an intermediate reward model. They are provided the same response pairs, which are sampled from Clarify SFT model. Specifically, we sample six candidate clarifying questions (one with $T=0.0$ and five with $T=1.0$) and form preference pairs among them. The only difference is how to assign reward. For Clarify DPO (Starling) model, we use the Starling-RM-7B-alpha reward model (Zhu et al., 2023) which was finetuned from Llama2-7B-Chat on over 3.8M single-turn preferences between the outputs from a variety of different LLMs. For Clarify DPO (Ours) model, we use the preference assigned by simulating double-turn preferences using our user-simulator model. We construct this model by finetuning the base LLM on all $(x,q,y_{i})\rightarrow a_{i}$ pairs.
Clarify-or-Direct Ans DPO (Ours)
@@
We first build Clarify-or-Direct Ans SFT model by training on the union of two datasets used to train Direct Ans SFT and Clarify SFT. We then further train this model over by double-turn preferences over clarifying question and direct-answer responses with DPO learning objective. We perform DPO training over pairwise comparisons from the same six candidate clarifying questions generated for our Clarify DPO models above and the single greedy sampled direct-answer response generated from our Direct Ans SFT model above.

### v2

@@ Abstract
Large language models (LLMs) must often respond to highly ambiguous user requests. In such cases, the LLM’s best response may be to ask a clarifying question to elicit more information. Existing LLMs often respond by presupposing a single interpretation of such ambiguous requests, frustrating users who intended a different interpretation. We speculate this is caused by current preference data labeling practice, where LLM responses are evaluated only on their prior contexts. To address this, we assign preference labels by simulating their expected outcomes in future turns. This allows LLMs to learn to ask clarifying questions when it can generate responses that are tailored to each user interpretation in future turns. On open-domain QA datasets with multiple annotations, we evaluate systems based on their ability to ask clarifying questions to recover each user’s interpretation and expected answer. We compare systems trained using our proposed preference labeling methods against standard methods, which assign preferences based on only prior context. Our method achieves a 5% improvement in F1 measured against the answer set from different interpretations of each query, showing the value of modeling future conversation turns. We further demonstrate that our method can be used to train models to judiciously determine when to ask clarifying questions, directly answering the question when clarification is unnecessary. In our experiments, we find that our method achieves a 3% improvement in accuracy of such judgments over existing methods.
@@ 1 Introduction
We propose an alternative method for annotating double-turn preferences over clarifying questions where annotators interact with the LLM by providing their responses to clarifying questions and observing the LLM’s subsequent responses. Each annotator then assigns preferences based on whether the completed interaction successfully fulfilled the request. When comparing multiple clarifying questions and direct answer responses, each annotator’s preferences are aggregated to identify the response that maximizes preference across all annotators. We depict our proposed double-turn preference annotation method and compare against standard methods in Figure 1.
To demonstrate the benefits of training LLMs to ask clarifying questions with double-turn preferences, we experiment on open-domain QA (Chen et al., 2017). We develop an automatic evaluation framework for evaluating clarifying questions that use simulated user interactions (Section 2). We evaluate systems on two axes: efficiency measured by the number of model turns and effectiveness measured by F1 score between the predicted answer set for each user and their expected gold answer.
Figure 1: Our interaction scenario and preference labeling schemes. We aim to build an LLM that can interact with user to generate the final answer set $R$, containing an answer for each user, for the input query $x$. In this example, we include two responses from state-of-the-art LLMs ([A] from GPT-4 and [B] from Gemini, full responses in Appendix A), which both presuppose an interpretation of the word football. We also include two clarifying responses ([C] and [D]) where [C] correctly disambiguates the two intended interpretations across all users. We depict two ways to assign preferences on LLM’s initial output, single-turn and our proposed double-turn.
Our experimental results show that training systems with double-turn preferences outperforms training with standard preferences annotation methods (Zhu et al., 2023), resulting in consistent 4-5% improvement in F1 score over three base LLMs (Dubey et al., 2024; Mesnard et al., 2024). We also demonstrate that double-turn preferences can be used to train systems that determine if clarification is needed, or if the user’s question can be answered without additional interaction.
@@ 2.1 Interaction Scenario
Setting / Goal: We assume an input query $x$, which can be potentially ambiguous, and a set of $k$ users $u_{1},...u_{k}$ who posed the input query $x$. Each $i$-th user $u_{i}$ has an disambiguated intent for query ($x_{i}$) and an expected gold answer ($y_{i}$). Our goal is to build a system $M$ that returns an answer set $R$ that can satisfy as many users as possible (by providing the expected answer $y_{i}$ for each user $u_{i}$) with the minimum number of interactions (measured by the number of model turns).
We will have a simple function $\phi(M(x))$ that returns True if $M(x)$ is a clarifying question, and False otherwise. We further assume an oracle user simulation model $\psi(x,r_{init},y_{i})=a_{i}$, that takes first turn input $x$, model’s initial output $M(x)$, and $i$-th user’s targeted answer $y_{i}$, that generates a response to the model’s initial response $M(x)$.
Dialogue Turns: In the initial turn, user poses the input query $x$. In the second turn, given the input query $x$, the system $M$ is to generate its initial response ($r_{\text{init}}$). Depending on whether $r_{\text{init}}$ is a question, it will trigger different dialogue trajectories.
@@
Existing QA models mostly generates an single answer to the initial input query $x$, even when the query is ambiguous. Such systems that predict an answer without first confirming the user’s intent may mislead users with incorrect answers and fail to serve the entirety of their user population. In this work, we explore methods for QA-with-Clarification, where models can ask the user a clarifying question $q$ and observe the user’s response $a$ before predicting an answer. A system can also sometimes poses a clarifying question and sometimes directly answer (Zhang & Choi, 2023a).
Evaluations: We evaluate the system in two axis: efficiency and effectiveness. To measure efficiency, we simply measure the average number of total turns from the system. For unambiguous inputs, asking clarifying question is unnecessary. If the system can satisfy users without incurring additional interaction, such system should be preferred.
To measure effectiveness, we use F1 between the model generated answer set ($R$) and the gold reference answer set ($\{y_{1},..y_{k}\})$. When evaluating QA-with-Clarification systems, we additionally enforce that answers must be correct for their corresponding clarifying answer. We hold QA-with-Clarifications to this higher standard, as it aligns with our goal of to developing systems that can identify different user interpretations of a query and accurately predict corresponding answers.
@@ 2.2 Evaluation Framework Implementation
Data: We will leverage existing open-domain QA datasets (Kwiatkowski et al., 2019; Min et al., 2020) where each query is paired with annotated answers from multiple annotators.
Identifying clarifying question $\phi(\cdot)$: We will use a very simple method, which takes the input string and check whether it starts with “Clarifying Question:”.
User Simulation Model $\psi$ for Turn 3: We use GPT-4 as our oracle user simulation model ($\psi$). We prompt GPT-4 with the input query $x$, clarifying question $q$, and target answer $y_{i}$ and task it with generating a corresponding clarifying answer $((x,q,y_{i})\rightarrow a_{i})$. We prompt GPT-4 to abstain from providing a clarifying answer $a_{i}$ if it judges that none exists. This happens when the clarifying question does not address the ambiguity in the query, hence we count the resulting target answer prediction $r^{i}_{\text{next}}$ as incorrect. We further ensure that the generated clarifying answers do not leak the target answer by removing instances where the target answer $y_{i}$ appears in the predicted clarifying answer $a_{i}$, treating such examples the same as abstains from GPT-4 (full prompt in Appendix C).
@@ 3 Fine-tuning LLMs to ask Clarifying Questions
We present our approach to build a model that can engage with users for a multi-turn interaction, ask clarifying question when it can satisfy more users in the future turns. Following the standard RLHF training pipeline (Ouyang et al., 2022), we first construct an instruction tuning dataset focusing on clarifying responses as desired outputs. We then use this constructed dataset for supervised finetuning (SFT), before performing preference learning between two possible responses. We first describe our method for constructing this instruction tuning dataset before introducing our annotation scheme for assigning preferences over clarifying questions.
@@ 3.1 Preference Data Generation from User-LLM Interactions
Our method for labeling preferences deviates from standard annotation methods (Bai et al., 2022; Wang et al., 2024) in two ways. First, we simulate an additional interaction with the user and derive reward from the LLM’s final response after the additional turn. Second, we expect multiple annotators for each example. By using multiple human annotators to identify a set of expected answers, we simulate each annotator’s interactions with the LLM and their preferences. While prior work has explored using multiple annotator judgments in single-turn preference datasets (Wang et al., 2024; Kopf et al., 2023), these works aggregate annotator judgments via majority choice to generate a single preference label per example. This practice, however, can lead to adverse outcomes where systems whose outputs are catered toward a single, majority accepted response (Fleisig et al., 2023; Santy et al., 2023). In contrast, we use preferences across multiple annotators without removing individual annotator judgments, helping us identify whether a clarifying question or direct-answer response successfully accommodates the range of different annotator’s interpretations.
@@
In our annotation scheme (depicted in Figure 2), annotators are provided an input query $x$ with several candidate clarifying question $q$ and direct-answer responses $\widehat{y}$. They respond to each clarifying question by providing the clarifying answer that corresponds to their interpretation. We then show annotators the model’s final output prediction, given each clarification. Annotators then assign their preferences over clarifying questions and direct-answer responses based on whether the final output matched their expected interpretation. We then aggregate these preferences to identify which response was preferred by the most annotators.
To prevent systems from asking unnecessary clarifying questions, ties between direct-answer and clarifying question responses are broken by favoring the direct-answer response. The final aggregated preferences can then be used in conjunction with standard RL methods from human preferences for training. While this annotation scheme can be applied to human annotators, in this work we use simulated user interactions for annotation, simulating different users’ interactions from existing datasets with multiple annotations (discussed below in Section 4).
@@
To simulate different user interpretations of the input query, we utilize datasets consisting of an input query $x$ and a set of correct answer outputs $\{y_{1},...,y_{k}\}$. For each query, we use the set of gold outputs to simulate $k$ different users, one for each answer in $\{a_{1},...,a_{k}\}$ which are identified by multiple human annotators (we discuss how these annotations are sourced in Section 4). We then simulate the behaviors of each of these users and their responses to a proposed clarifying question $q$ using a trained user-simulator model, which takes as input query $x$, clarifying question $q$, and the user’s annotated answer $y_{i}$ and is trained to predict the user’s clarifying answer $(x,q,y_{i})\rightarrow a_{i}$. While we use prompted GPT-4 to simulate clarifying answers from users for evaluation, during training we use this trained user-simulator model to reduce costs and to ensure that our systems do not overfit to a particular user-simulator method.
@@ 4 Experimental Settings
We use Llama2-7b (Touvron et al., 2023), Gemma-7b (Mesnard et al., 2024), and Llama3-8b (Dubey et al., 2024) as our base LLMs. We forego using the instruction-tuned models due to data leakage concerns, as such systems are finetuned on the entirety of NQ-Open. For training and inference, we use 8-bit quantization (Dettmers et al., 2022) with LoRA (Hu et al., 2022; Dettmers et al., 2024) (training details in Appendix C).
@@
  • AmbigQA additionally annotates a subset of NQ-Open whether the question is ambiguous. If the input is ambiguous, annotators then provide the additional answers to other possible interpretations of the query. This process identifies about 56% of all queries in NQ-Open as ambiguous, and recovers 2.6x more answers to ambiguous questions.
@@
  • Answers from Model Predictions: We build feasible answer sets $Y_{\text{model}}$ for queries in the NQ-Open training set using answer candidates proposed by a base LLM. Prior work has demonstrated that base LLMs are well calibrated in open-domain QA tasks with few-shot prompting (Kadavath et al., 2022; Cole et al., 2023). Drawing on these observations, we construct a feasible answer set $Y_{\text{model}}$ by sampling 5-shot predictions from Llama-2-7b (Touvron et al., 2023). For each query, we sample 5-shot prompt and generate its greedy answer with temperature $T=0.0$ and sample an answer with $T=1.0$. We repeat the process 10 times per query and remove repeated answers to generate a set of model-identified feasible answers. We further filter examples where none of the feasible answers match any annotated answers from NQ-Open.
@@ 4.1 Compared Systems
Figure 3 summarizes all models used in our experiments. We have two types of models, model that will be used at the second turn (which we refer as Turn 2 model) and model that will be used at fourth turn (which we refer as Turn 4 model and notate as “Ans-After-Clarify” model in the figure). The model that will be used at the second turn can ask clarification question or directly answer the question. All models that are capable of asking clarification questions are paired with the same Turn 4 model, which is trained over all generated SFT examples. We use a separate model for performing this latter step to ensure that differences in performance only reflect the value of the clarifying question, and explore training joint models for both steps in Section 5. We describe each of these below.
@@
Direct Ans SFT This model is trained to predict the answer given the input query $(x)\rightarrow y$. We fine-tune base LLM with next-token prediction loss on the full set of question answer pairs (i.e., each $(x,y_{i})$) in our generated SFT training sets. We use this model under two inference settings: (1) Greedy where we take the single greedy answer prediction and (2) Sampled where we sample a set predicted answers from the LLM. We sample 20 answers from the model and select the top $k$ most frequent answers, where $k$ matches the gold number of annotated answers. [^1] In rare cases ($<2\%$), there are less than $k$ unique answers within the 20 samples.
@@
Clarify SFT This model is trained to ask clarifying question to input query. We finetune base LLMs on our SFT datasets constructed with both human annotations and model predicted answers. In Section 5, we provide ablation experiments training over each dataset to compare their utility.
Clarify DPO (RM, Likelihood, Match) These models are each fine-tuned from Clarify SFT model with using DPO (Rafailov et al., 2023), a method for optimizing LLMs on preference data. It uses the following loss for updating the policy LLM $\pi_{\theta}$ based on the base reference policy $\pi_{ref}$, prompt $h$, and preferred/rejected responses $(r_{p},r_{r})$:
$L(\pi_{\theta},\pi_{ref},r_{p},r_{r},h)=\log\sigma(\beta\log\frac{\pi_{\theta}(r_{p}|h)}{\pi_{ref}(r_{p}|h)}-\beta\log\frac{\pi_{\theta}(r_{r}|h)}{\pi_{ref}(r_{r}|h)})$
where $\beta$ is a hyperparameter and $\pi(r|h)$ is the predicted probability of generating $r$ given the prompt $h$. These models are provided the same response pairs sampled from Clarify SFT model, [^2] We sample six candidate clarifying questions (one with $T=0.0$ and five with $T=1.0$) and form preference pairs among them. only differing in how reward are assigned.
  • RM: This uses off-the-shelf reward model to rank clarifying questions. We use the Starling-RM-7B-alpha reward model (Zhu et al., 2023) which was finetuned on over 3.8M single-turn preferences between the outputs from a variety of LLMs.
  • Likelihood: In this method, clarifying questions are ranked by the summed likelihoods of generating each gold answer after observing the simulated users’ clarifying answers. Using likelihood to score questions has been proposed in prior work studying clarifying question generation to identify different user personas (Andukuri et al., 2024).
  • Match: We rank clarifying questions by whether the predicted answer matches each user’s gold answer after observing each simulated user’s clarifying answer, averaged over all simulated users.
@@
Clarify-or-Direct Ans SFT / DPO We train a Clarify-or-Direct Ans SFT model on the union of all data used to train our Clarify SFT and Direct-Ans SFT models above. To train our Clarify-or-Direct Ans DPO model, wefurther train our Clarify-or-Direct Ans SFT model above using double-turn preferences over clarifying question and direct-answer responses with the DPO learning objective. To rank responses, we use answer matching as depicted in Figure 2. We perform DPO training over pairwise comparisons from the same six candidate clarifying questions generated for our Clarify DPO models above and the single greedy sampled direct-answer response generated from our Direct Ans SFT model above.